# Isometries of spacetimes without observer horizons

### v2

@@ Abstract.
We thank the organizers of the conference “Contact and Lorentzian Geometry II” at the Ruhr-Universität Bochum, where this collaboration started.
@@ 1. Introduction
On compact spacetimes, it is known that the causal relation cannot be anti-symmetric, because there always exist closed causal curves. Because of this, spacetimes considered in general relativity are generally non-compact, since closed causal curves would imply the possibility of time-travel, which is considered unphysical. In the present paper, we will see that constraints on the causal relation lead to isometries being quite rigid, because they have to preserve the global causal structure. This contrasts with the defining property of a isometry, i.e. preserving the metric tensor, which is merely local.
In addition to causality (i.e. antisymmetry of $\leq$), we assume the “no future observer horizons” condition (NFOH), which prevents the lightcones from becoming excessively narrow towards infinity. Equivalently, it ensures that every observer is able to receive signals from every point in spacetime (after waiting long enough). The precise definition is that $I^{+}(\gamma)=M$ for every future-inextendible causal curve $\gamma$. This condition has been extensively studied in the context of the causal boundary and Bartnik’s splitting conjecture [13]. As pointed out by Penrose [31], the combination of causality and the NFOH implies another standard causality assumption, namely global hyperbolicity with compact Cauchy surfaces. It has been shown by various authors that spacetimes with certain kinds of “time-translation like” symmetry must satisfy the NFOH [8, 11, 16]. Here, we turn the situation around, and study how the NFOH restricts the possible symmetries of the spacetime.
@@
There exists a Cauchy temporal function $\tau\colon M\to\mathbb{R}$ such that the action of $\operatorname{Isom}^{\uparrow}(M,g)$ on $M$ preserves the gradient one-form $d\tau$.
@@
Properness of the action (which we recall later on) means essentially the action has a non-chaotic dynamics, so that all orbits are closed, and even more, the orbit space is Hausdorff. Corollary 1.3 shows that, when the NFOH is satisfied, the splitting of the spacetime into a product of time and space via a Cauchy temporal function (due to Geroch [17] and Bernal and Sánchez [5]) induces a splitting of the isometry group. The NFOH is essential here; a counterexample is given by de Sitter spacetime, which has compact Cauchy surfaces, but does not satisfy any of our conclusions. See also Müller [27] for prior work on temporal functions invariant under a Lie group action, where the spacetime is only assumed globally hyperbolic, but the Lie group is assumed compact.
@@
The paper is structured as follows: In Section 2 we discuss preliminaries on both Lie group actions and Lorentzian geometry. The goal is both to make the paper more accessible, and to prove certain basic results that are not found in the literature in that precise form. Section 3 contains the proofs of the results stated in this introduction (restated there with some added detail), and many examples. In Section 4, we have included some additional results and examples. In particular, we prove uniform bi-Lipschitz continuity of certain subgroups of isometries, and give an application to spacetimes equipped with a regular cosmological time function.
@@ 2.1. Lie groups
Now let $G$ be a Lie group. Its Lie algebra $\mathfrak{g}$ is the vector space of smooth left-invariant vector fields on $G$, equipped with the usual Lie bracket $\cdot$$\cdot$]. There is a canonical isomorphism of vector spaces $\mathfrak{g}\to T_{e}G$ between the Lie algebra and the tangent space at the identity $e\in G$. The exponential map $\exp\colon\mathfrak{g}\to G$ sends a vector $X\in\mathfrak{g}$ to $\exp(X):=\Phi(1)\in G$, where $\Phi\colon\mathbb{R}\to G$ is the unique one-parameter subgroup with $\Phi^{\prime}(0)=X$. A word of caution is appropriate here: We will also use $\exp$ to denote the exponential map on a semi-Riemannian manifold (defined through geodesics); it should be clear from context which one is meant.
@@ 2.2. Group actions
Theorem 2.3 below will be a central tool in many of our proofs. A weaker version of it can be found in the book of Kobayashi [21, Chap. II.1] (see also the lecture notes by Ballmann [3], which we follow), where it is used to prove that the isometry group of a (semi-)Riemannian manifold is a Lie group. Before we start, we need the following well-known lemma.
@@
We follow [3, Sec. 3]. Let $\pi\colon O(M)\to M$ denote the canonical projection. Our first goal is to split the tangent bundle $TO(M)$ into a vertical and a horizontal part, $TO(M)=V\oplus H$. Here $V=\ker D\pi$, and we want $D\pi\colon H_{F_{p}}\to T_{p}M$ to be an isomorphism of vector spaces. While $V$ is uniquely defined, $H$ is not, and our task is to find a such an $H$ that is compatible with our geometric structure (i.e. with the metric $g$).
@@
Let $\Phi$ be the parallelization of $O(M)$ obtained in Lemma 2.2. By [3, Prop. 1.5], the action of the automorphism group $\operatorname{Aut}(\Phi)$ is free, and the orbit map of any given $F_{p}\in O(M)$,
@@ 2.3. Lorentzian geometry
Moreover, any two Cauchy surfaces are diffeomorphic, and every Cauchy temporal function $\tau$ gives rise to a diffeomorphism $M\to\mathbb{R}\times\Sigma$ such that $\tau$ coincides with the canonical projection.
@@
It is equivalent to require the same conditions only on timelike curves. Moreover, the NFOH (NPOH) is equivalent to the future (past) causal boundary consisting of a single point. We formulate most results for the NFOH, but there is always an analogous version for the NPOH. We first prove that the NFOH upgrades one of the lowest levels on the causal ladder, causality, to the highest level, global hyperbolicity (see also [31, p. 624] and [9, Sec. 9]).
@@
First, we prove that $(M,g)$ is globally hyperbolic. Since $(M,g)$ is assumed causal, we only need to prove that the causal diamonds $J(p,q):=J^{+}(p)\cap J^{-}(q)$ are compact (see Theorem 2.5). Suppose that there are $p,q\in M$ with $J(p,q)$ non-compact. Then, there is a sequence $x_{i}\in J(p,q)$ with no converging subsequences. Since $x_{i}\in J^{+}(p)$, there is a corresponding sequence of causal curves $\gamma_{i}$ from $p$ to $x_{i}$. By the limit curve theorem [24, Thm. 2.53], there is a subsequence $\gamma_{i_{k}}$ that converges to a limit curve $\gamma$ starting from $p$, and one of the following holds:
  1. (i)
    $\gamma$ is future-inextendible.
  2. (ii)
    $\gamma$ has $\lim_{k\to\infty}x_{i_{k}}$ as its future-endpoint.
The second option is impossible, given that $x_{i}$ had no converging subsequences. Therefore, $\gamma$ is future-inextendible. By the NFOH, $I^{-}(\gamma)=M$. In particular, $q\in I^{-}(\gamma)$. Thus, there is a point $\gamma(s_{0})$ such that $q\ll\gamma(s_{0})$. By openness of the chronological relation, for $i$ large enough, $q\ll\gamma_{i}(s_{0})$. But because $\gamma_{i}\subset I^{-}(q)$, this contradicts causality.
Having established that $(M,g)$ is globally hyperbolic, it follows from Theorem 2.5 that there is a Cauchy surface $\Sigma\subset M$. It follows directly from [33, Lem. 2] that $\Sigma$ is compact (note that the NFOH implies that the future causal boundary is a singleton, and hence it is spacelike). ∎
@@
Because the level sets $\Sigma_{t}$ are achronal, by the NOH and Lemma 2.9 we can find a pair of totally timelike connected Cauchy surfaces such that $\Sigma_{t}$ lies in between them. Then $\Sigma_{t}$ is a Cauchy surface by [12, Cor. 2]. ∎
@@ 3.1. Proofs of the main results
By Proposition 2.1, the action is proper if and only if for every pair of sequences $p_{i}\in M$ and $\phi_{i}\in\operatorname{Isom}^{\uparrow}(M,g)$ such that $p_{i}$ and $\phi_{i}(p_{i})$ converge in $M$, a subsequence of $\phi_{i}$ converges in $\operatorname{Isom}^{\uparrow}(M,g)$. By Theorem 2.3, we know that this will be the case if there is a converging sequence of frames $F_{i}$ at $p_{i}$, for which a subsequence of $D\phi_{i}(F_{i})$ converges in $O(M)$.
@@
Suppose that we have a converging sequence of frames $F_{i}$ at $p_{i}\to p$ such that $D\phi_{i}(F_{i})$ has converging basepoints but becomes degenerate in the limit. This can only happen if there is a sequence of spacelike vectors $v_{i}$ at $p_{i}$ that converges to a spacelike $v\neq 0$, such that $w_{i}:=c_{i}D\phi_{i}(v_{i})$ converges to a future-directed null vector $w$ at $q:=\lim_{i\to\infty}\phi_{i}(p_{i})$. Here $c_{i}$ is a sequence of positive numbers, which necessarily tends to 0, chosen so that $w_{i}$ converges (for instance, the $c_{i}$’s can be normalization constants with respect to an auxiliary Riemannian metric). This situation is depicted in Figure 1.
By Proposition 2.7, $(M,g)$ is globally hyperbolic, and hence strongly causal. Thus, there is an $\varepsilon>0$ such that for all $t\in(0,\varepsilon]$ and all $i\in\mathbb{N}$, $p_{i}\not\ll\exp(tv_{i})$. We may also multiply all the $w_{i}$ by the same constant $c^{\prime}$, so that $w^{\prime}_{i}:=c^{\prime}w_{i}\to w^{\prime}:=c^{\prime}w$. Since $s\mapsto\exp(sw)$ is an inextendible causal curve, the NFOH implies that for $c^{\prime}$ large enough, $q\ll\exp(w^{\prime})$. On the one hand, by continuity of $\exp$ and openness of $\ll$, this implies that $\phi_{i}(p_{i})\ll\exp(w^{\prime}_{i})$ for all large enough $i$. On the other hand, because the $\phi_{i}$ are isometries,
@@
    There exists a Cauchy temporal function $\tau\colon M\to\mathbb{R}$ such that the action of $G$ on $M$ preserves the gradient one-form $d\tau$.
@@
is compact in $G$ (by propreness of the action), and hence the integral (5) is finite.
@@
By Lemma 2.9, there are real numbers $a<b<c<d<e$ such that the corresponding $\hat{\tau}$ level sets $\Sigma_{a}$, $\Sigma_{b}$, $\Sigma_{c}$, $\Sigma_{d}$, and $\Sigma_{e}$ are pairwise totally timelike connected. Consider the compact set $A:=\hat{\tau}^{-1}((a,e))=I^{+}(\Sigma_{a})\cap I^{-}(\Sigma_{e})$ and a point $p\in\Sigma_{c}\subset A$. By propreness of the action, if $G$ is non-compact, then there must be $\phi\in G$ such that $\phi(p)\not\in A$. Since $\Sigma_{c}$ is a Cauchy surface, so is $S_{1}:=\phi(\Sigma_{c})$, and it follows by achronality that $S_{1}\cap\hat{\tau}^{-1}((b,d))=\emptyset$. Therefore $\Sigma_{c}$ and $S_{1}$ are totally timelike connected (or $S_{1}$ and $\Sigma_{c}$, but then we may replace $\phi$ by $\phi^{-1}$, so there is no loss in generality). Moreover, we recursively construct a family of Cauchy surfaces $S_{0}:=\Sigma_{c}$, $S_{i}:=\phi^{i}(S_{0})$ for all $i\in\mathbb{Z}$ with $S_{i+1}\subset I^{+}(S_{i})$. We claim that every $G$-orbit contains a representative in the compact set
@@
$L(\gamma)=\sum_{i=0}^{\infty}L(\gamma_{i})>\sum_{i=0}^{\infty}D_{i}=\infty.$
@@
We end this section by proving the last part of Corollary 1.3, i.e. that when $\operatorname{Isom}^{\uparrow}(M,g)\cong\mathbb{R}\ltimes N$, then the connected component of the identity splits as a direct (and not just semi-direct) product. The relevant lemma about Lie groups below can also be found in [19, Lem. 5.3]. Its proof requires connectedness, for two reasons: to have a surjective exponential map, allowing us to argue with the Lie algebra, and to be able to pass to the universal covering group.
@@
We first deal with the case $N=\mathbb{T}^{k}\times S$ for $S$ a compact, connected, and semi-simple Lie group. Since $L\cong\mathbb{R}$, we may view $\Psi\colon L\to\operatorname{Aut}(N)$ as a one-parameter subgroup $s\mapsto\Psi(s)$. Note that $\operatorname{Aut}(N)=\operatorname{Aut}(\mathbb{T}^{k})\times\operatorname{Aut}(S)$, because every automorphism of $N$ must preserve the identity component of the center $Z^{0}(N)=\mathbb{T}^{k}$. Since $\operatorname{Aut}(\mathbb{T}^{k})=\operatorname{GL}(k,\mathbb{Z})$ is discrete, $\Psi(s)\in\operatorname{Aut}^{0}(S)$ for all $s\in\mathbb{R}$ (extended trivially to $N$). Note that because $S$ is compact and connected, by [20, Cor 4.48], the exponential map $\exp\colon\mathfrak{s}\to S$ is surjective. Therefore, the automorphism $\Psi(s)\in\operatorname{Aut}(S)$ is uniquely determined by $d\Psi(s)\in\operatorname{Aut}(\mathfrak{s})$. Then $s\mapsto d\Psi(s)$ is a one parameter subgroup of $\operatorname{Aut}(\mathfrak{s})$. By [20, Chap. I.14], the Lie algebra of $\operatorname{Aut}(\mathfrak{s})$ is $\operatorname{Der}(\mathfrak{s})=\operatorname{ad}(\mathfrak{s})$. Therefore, there exists a $V\in\mathfrak{s}$ such that
@@
It follows by the above and [20, Prop. 1.91] that $d\Psi(s)=e^{\operatorname{ad}(sV)}=\operatorname{Ad}(\exp(sV))$, where $e^{(\cdot)}\colon\operatorname{Der}(\mathfrak{s})\to\operatorname{Aut}(\mathfrak{s})$ is the exponential map, and hence
@@
It remains to deal with the case of general $N$. We do this by passing to a covering of the form dealt with in the previous case. By [20, Prop. 1.124] and [20, Cor. 4.25], the Lie algebra $\mathfrak{g}$ of $G$ is of the form
@@
where $\psi$ is the differential of $\Psi$, and $\mathfrak{s}$ is a compact semi-simple Lie algebra. Therefore, there is a compact, simply connected, semi-simple Lie group $S$ with Lie algebra $\mathfrak{s}$. Then $\mathbb{R}^{k}\times S$ has Lie algebra $\mathbb{R}^{k}\oplus\mathfrak{s}$. By [20, Thm. 1.125], there exists a unique map $\Phi\colon L\to\operatorname{Aut}(\mathbb{R}^{k}\times S)$ such that

### v3

@@ Abstract.
We thank the organizers of the conference “Contact and Lorentzian Geometry II” at the Ruhr-Universität Bochum, where this collaboration started. LGH thanks Eric Ling for interesting conversations about cosmology. This work was supported in part by the Austrian Science Fund (FWF) [Grant DOI 10.55776/EFP6]. For open access purposes, the authors have applied a $C$-BY public copyright license to any author accepted manuscript version arising from this submission.
@@ 1. Introduction
On compact spacetimes, it is known that the causal relation cannot be anti-symmetric, because there always exist closed causal curves. Because of this, spacetimes considered in general relativity are generally non-compact, since closed causal curves would imply the possibility of time-travel, which is considered unphysical. In the present paper, we will see that constraints on the causal relation lead to isometries being quite rigid, because they have to preserve the global causal structure. This contrasts with the defining property of an isometry, i.e. preserving the metric tensor, which is merely local.
In addition to causality (i.e. antisymmetry of $\leq$), we assume the “no future observer horizons” condition (NFOH), which ensures that every observer is able to receive signals from every point in spacetime (after waiting long enough). The precise definition is that $I^{-}(\gamma)=M$ for every future-inextendible causal curve $\gamma$. This condition has been extensively studied in the context of the causal boundary and Bartnik’s splitting conjecture [13]. As pointed out by Penrose [34], the combination of causality and the NFOH implies another standard causality assumption, namely global hyperbolicity with compact Cauchy surfaces. It has been shown by various authors that spacetimes with certain kinds of “time-translation like” symmetry must satisfy the NFOH [9, 15, 19]. Here, we turn the situation around, and study how the NFOH restricts the possible symmetries of the spacetime.
@@
There exists a Cauchy temporal function $\tau\colon M\to\mathbb{R}$ such that the action of $\operatorname{Isom}^{\uparrow}(M,g)$ on $M$ preserves the differential $d\tau$.
@@
Properness of the action (which we recall later on) means essentially the action has a non-chaotic dynamics, so that all orbits are topologically closed, and even more, the orbit space is Hausdorff. Corollary 1.3 shows that, when the NFOH is satisfied, the splitting of the spacetime into a product of time and space via a Cauchy temporal function (due to Geroch [20] and Bernal and Sánchez [6]) induces a splitting of the isometry group. The NFOH is essential here; a counterexample is given by de Sitter spacetime, which has compact Cauchy surfaces, but does not satisfy any of our conclusions. See also Müller [30] for prior work on temporal functions invariant under a Lie group action, where the spacetime is only assumed globally hyperbolic, but the Lie group is assumed compact.
@@
The paper is structured as follows: In Section 2 we discuss preliminaries on Lie group actions, Lorentzian geometry, and the physical context. The goal is both to make the paper more accessible, and to prove certain basic results that are not found in the literature in that precise form. Section 3 contains the proofs of the results stated in this introduction (restated there with some added detail), and many examples. In Section 4, we have included some additional results and examples. In particular, we prove uniform bi-Lipschitz continuity of certain subgroups of isometries, and give an application to spacetimes equipped with a regular cosmological time function.
@@ 2.1. Lie groups
Now let $G$ be a Lie group. Its Lie algebra $\mathfrak{g}$ is the vector space of smooth left-invariant vector fields on $G$, equipped with the usual Lie bracket $\cdot$$\cdot$]. There is a canonical isomorphism of vector spaces $\mathfrak{g}\to T_{e}G$ between the Lie algebra and the tangent space at the identity $e\in G$. The exponential map $\exp\colon\mathfrak{g}\to G$ sends a vector $X\in\mathfrak{g}$ to $\exp(X):=\varphi(1)\in G$, where $\varphi\colon\mathbb{R}\to G$ is the unique one-parameter subgroup with $\varphi^{\prime}(0)=X$. A word of caution is appropriate here: We will also use $\exp$ to denote the exponential map on a semi-Riemannian manifold (defined through geodesics); it should be clear from context which one is meant.
@@ 2.2. Group actions
Theorem 2.3 below, telling us that the isometry group acts properly on $O(M)$, will be a central tool in many of our proofs. We are not aware of a reference where it appears in this exact form, but similar ideas can be found in the book of Kobayashi [24, Chap. II.1] and, with more details, in the lecture notes by Ballmann [4]. In those references, it is proven that the isometry group of a (semi-)Riemannian manifold is a Lie group. Before we start, we need the following well-known lemma.
@@
We follow [4, Sec. 3]. Let $\pi\colon O(M)\to M$ denote the canonical projection. Our first goal is to split the tangent bundle $TO(M)$ into a vertical and a horizontal part, $TO(M)=V\oplus H$. Here $V=\ker D\pi$, and we want $D\pi\colon H_{F_{p}}\to T_{p}M$ to be an isomorphism of vector spaces. While $V$ is uniquely defined, $H$ is not, and our task is to find a such an $H$ that is compatible with our geometric structure (i.e. with the metric $g$).
@@
Let $\Phi$ be the parallelization of $O(M)$ obtained in Lemma 2.2. By [4, Prop. 1.5], the action of the automorphism group $\operatorname{Aut}(\Phi)$ is free, and the orbit map of any given $F_{p}\in O(M)$,
@@ 2.3. Lorentzian geometry
Moreover, any two Cauchy surfaces are diffeomorphic, and every Cauchy temporal function $\tau$ gives rise to a diffeomorphism $M\to\mathbb{R}\times\Sigma$ such that $\tau$ coincides with the first projection.
@@
It is equivalent to require the same conditions only on timelike curves [11, Prop. 3.32]. Moreover, the NFOH (NPOH) is equivalent to the future (past) causal boundary consisting of a single point. We formulate most results for the NFOH, but there is always an analogous version for the NPOH. We first prove that the NFOH upgrades one of the lowest levels on the causal ladder, causality, to the highest level, global hyperbolicity (see also [34, p. 624] and [10, Sec. 9]).
@@
The NFOH is equivalent to the absence of future lightlike rays, i.e. future inextendible achronal causal curves. Clearly, if $\gamma$ is future-inextendible, then $I^{-}(\gamma)=M$ and $\gamma$ cannot be achronal. Conversely, if $I^{-}(\sigma)\neq M$ for some future-inextendible causal $\sigma$, then $\partial I^{-}(\sigma)$ is non-empty and is ruled by future inextendible lightlike rays. Minguzzi proved that if $(M,g)$ is chronological and has no future lightlike rays, then it is globally hyperbolic [27, Thm. 8], settling the first part of the proposition. Then Theorem 2.5 implies there is a Cauchy surface $\Sigma\subset M$. A result of Wald and Yip [39, Lem. 2] states that for every future inextendible causal curves $\gamma$, $\Sigma\cap I^{-}(\gamma)$ is compact, which in our case means that $\Sigma\cap M=\Sigma$ is compact (note that the NFOH implies that the future causal boundary is a singleton, and hence it is spacelike, so the assumptions of [39, Lem. 2] are met). ∎
@@
By Proposition 2.7, $(M,g)$ is globally hyperbolic, and hence admits a Cauchy temporal function $\tilde{\tau}$. Let $\Sigma_{t}$ be any level set of $\tau$, and $p\in\Sigma_{t}$. By Lemma 2.9, we can find level sets $\tilde{\Sigma}_{t_{i}}$ of $\tilde{\tau}$ for values $t_{1}<t_{2}=\tilde{\tau}(p)<t_{3}$ such that $\tilde{\Sigma}_{t_{i}}$ and $\tilde{\Sigma}_{t_{i+1}}$ are totally timelike connected. Because $\Sigma_{t}$ is achronal and intersects $\tilde{\Sigma}_{t_{2}}$, it cannot intersect $I^{-}(\tilde{\Sigma}_{t_{1}})$ or $I^{+}(\tilde{\Sigma}_{t_{3}})$. Thus $\Sigma_{t}\subset J^{+}(\tilde{\Sigma}_{t_{1}})\cap J^{-}(\tilde{\Sigma}_{t_{3}})$ and by [16, Cor. 2], $\Sigma_{t}$ is a Cauchy surface. ∎
Note that if $(M,g)$ only satisfies the NFOH, the statement is false in general. One can adapt the above proof, however, to show that when the NFOH is satisfied, given any time function $\tau$, all level sets corresponding to values larger than some constant are Cauchy surfaces. For example, on the spacetime $(0,\infty)\times S^{1}$ with product metric $dt^{2}+d\theta^{2}$, the level sets of the time function $\tau(t,\theta)=t+1/2\sin(\theta)$ are not Cauchy for values of $\tau$ smaller than $1/2$.
2.4. Physical context
Let us conclude this section with a physics oriented discussion of the NFOH and our results. In fact, it is more reasonable to consider the NPOH instead. By Proposition 2.7, our spacetimes will be globally hyperbolic with compact Cauchy surfaces. This places us in a cosmological context, since spacetimes modelling isolated systems (rather than the whole Universe) are usually asymptotically flat, and hence have non-compact Cauchy surfaces. Even in cosmology, the observed spatial flatness of the Universe has popularized FLRW spacetimes with flat slices (isometric to $\mathbb{R}^{n}$, hence non-compact). Recall that an FLRW spacetime is a warped product $((0,\infty)\times\Sigma,-dt^{2}+a^{2}(t)h)$, where $(\Sigma,h)$ is one of the Riemannian model spaces of constant curvature, and $a(t)$ is called the scale factor.
There are, however, reasons not to discard compact Cauchy surfaces altogether. Recently, the notion that spatial isotropy leads one to consider only FLRW spacetimes has been challenged [37, 3]. Specifically, Sánchez has introduced cosmological models which are foliated by spacelike slices of constant curvature, but which are not FLRW spacetimes, and where the sign of the curvature and the topology of the slices can change. In particular, there can be flat (non-compact, non-Cauchy) slices, even if the spacetime is globally hyperbolic with compact Cauchy surfaces.
Moving on, recall that for FLRW spacetimes, there is the no particle horizons condition
(4) $\int_{0}^{1}\frac{1}{a(t)}dt=\infty.$
In the case of compact spatial slices $\Sigma$, this condition is equivalent to the NPOH (see [38, Prop. 3.6] for a proof, which is valid for more general warped products, as it is not used that the slices are simply connected and of constant curvature). Current inflationary cosmology stipulates that there are no particle horizons, or at least that they are very large (i.e. the integral (4) is very large) [40, Chap. 4]. This is because the observed isotropy of the cosmic microwave background is explained by thermalization, which requires different spatial regions to have been in causal contact in the early universe. The characterization of the NPOH given by Lemma 2.9 provides a very strong version of this: any two points in space are causally connected in their past.
Let us use this opportunity to also discuss the intuitive meaning of the NFOH. One can interpret it as requiring that the spacetime has compact Cauchy surfaces (as is established by Proposition 2.7) and that additionally, the lightcones do not become excessively narrow towards future infinity (analogously for past infinity when considering the NPOH). Choosing a Cauchy splitting of the metric
$M=\mathbb{R}\times\Sigma,\quad g=-\beta dt^{2}+h=\beta(-dt^{2}+h/\beta),$
a sufficient condition for NFOH is
$g_{t}/\beta\leq h(t)g_{0}/(\beta|_{t=0})$
where $h(t)$ cannot diverge too fast when $t\to+\infty$ (this meaning that the cones are not narrowing too fast towards the future). This can be made precise in terms of the future arrival time function [36, Prop. 4.3]. Using this language, Lemma 2.9 can also be proven via continuity of the future arrival time function [36, Lem. 4.1].
@@ 3.1. Proofs of the main results
By Proposition 2.1, the action is proper if and only if for every pair of sequences $p_{i}\in M$ and $\phi_{i}\in\operatorname{Isom}^{\uparrow}(M,g)$ such that $p_{i}$ and $q_{i}:=\phi_{i}(p_{i})$ converge in $M$, a subsequence of $\phi_{i}$ converges in $\operatorname{Isom}^{\uparrow}(M,g)$. By Theorem 2.3, we know that this will be the case if there is a converging sequence of frames $F_{i}$ at $p_{i}$, for which a subsequence of $D\phi_{i}(F_{i})$ converges in $O(M)$.
@@
Suppose that we have a converging sequence of frames $F_{i}$ at $p_{i}\to p$ such that $D\phi_{i}(F_{i})$ has converging basepoints $q_{i}$ but becomes degenerate in the limit. This can only happen if there is a sequence of spacelike vectors $v_{i}$ at $p_{i}$ that converges to a spacelike $v\neq 0$, such that $w_{i}:=c_{i}D\phi_{i}(v_{i})$ converges to a future-directed null vector $w$ at $q:=\lim_{i\to\infty}\phi_{i}(p_{i})$. Here $c_{i}$ is a sequence of positive numbers, which necessarily tends to 0, chosen so that $w_{i}$ converges (for instance, the $c_{i}$’s can be normalization constants with respect to an auxiliary Riemannian metric). This situation is depicted in Figure 1.
By Proposition 2.7, $(M,g)$ is globally hyperbolic, so we may choose a convex normal neighborhood $U$ of $p$. Then, there is $\varepsilon>0$ small enough such that for all $t\in(0,\varepsilon]$ and all $i\in\mathbb{N}$, we have $p_{i},\exp(tv_{i})\in U$ and $p_{i}\not\ll\exp(tv_{i})$. We may also multiply all the $w_{i}$ by the same constant $c^{\prime}$, so that $w^{\prime}_{i}:=c^{\prime}w_{i}\to w^{\prime}:=c^{\prime}w$. Since $s\mapsto\exp(sw)$ is an inextendible causal curve, the NFOH implies that for $c^{\prime}$ large enough, $q\ll\exp(w^{\prime})$. On the one hand, by continuity of $\exp$ and openness of $\ll$, this implies that $\phi_{i}(p_{i})\ll\exp(w^{\prime}_{i})$ for all large enough $i$. On the other hand, because the $\phi_{i}$ are isometries,
@@
    There exists a Cauchy temporal function $\tau\colon M\to\mathbb{R}$ such that the action of $G$ on $M$ preserves the differential $d\tau$.
@@
is compact in $G$ (by properness of the action), and hence the integral (6) is finite.
@@
By Lemma 2.9, there are real numbers $a<b<c<d<e$ such that the corresponding $\hat{\tau}$ level sets $\Sigma_{a}$, $\Sigma_{b}$, $\Sigma_{c}$, $\Sigma_{d}$, and $\Sigma_{e}$ are pairwise totally timelike connected. Consider the compact set $A:=\hat{\tau}^{-1}((a,e))=I^{+}(\Sigma_{a})\cap I^{-}(\Sigma_{e})$ and a point $p\in\Sigma_{c}\subset A$. By properness of the action, if $G$ is non-compact, then there must be $\phi\in G$ such that $\phi(p)\not\in A$. Since $\Sigma_{c}$ is a Cauchy surface, so is $S_{1}:=\phi(\Sigma_{c})$, and it follows by achronality that $S_{1}\cap\hat{\tau}^{-1}((b,d))=\emptyset$. Therefore $\Sigma_{c}$ and $S_{1}$ are totally timelike connected (or $S_{1}$ and $\Sigma_{c}$, but then we may replace $\phi$ by $\phi^{-1}$, so there is no loss in generality). Moreover, we recursively construct a family of Cauchy surfaces $S_{0}:=\Sigma_{c}$, $S_{i}:=\phi^{i}(S_{0})$ for all $i\in\mathbb{Z}$ with $S_{i+1}\subset I^{+}(S_{i})$. We claim that every $G$-orbit contains a representative in the compact set
@@
(8) $L(\gamma)=\sum_{i=0}^{\infty}L(\gamma_{i})>\sum_{i=0}^{\infty}D_{i}=\infty.$
@@
Remark 3.3 (Comments on the proof of Theorem 3.2).
In the case of compact $G$, there cannot be an isometry $\phi$ acting on $\mathbb{R}$ by a non-trivial translation (otherwise, its powers $\phi^{k}$ would form a sequence with no convergent subsequence). Therefore $\tau$ itself, and not only $d\tau$, is invariant. A similar construction of an invariant temporal function was given by Müller in [30]. In the case of non-compact $G$, it follows from the proof that there is an isometry $\phi$ that not only acts by a non-trivial time translation, but even has the property that $p\ll\phi(p)$ for every $p\in M$. In our follow-up work [17], conformal transformations with this property are called escaping, and a similar argument with a fundamental domain $K$ is performed. In the present paper, however, we exploit the fact that we are working with isometries in equation (8), and this allows us to show that when there is an escaping isometry $\phi$, the NFOH is automatically promoted to the NOH.
@@
We end this section by proving the last part of Corollary 1.3, i.e. that when $\operatorname{Isom}^{\uparrow}(M,g)\cong\mathbb{R}\ltimes N$, then the connected component of the identity splits as a direct (and not just semi-direct) product. The relevant lemma about Lie groups below can also be found in [22, Lem. 5.3]. Its proof requires connectedness, for two reasons: to have a surjective exponential map, allowing us to argue with the Lie algebra, and to be able to pass to the universal covering group.
@@
We first deal with the case $N=\mathbb{T}^{k}\times S$ for $S$ a compact, connected, and semi-simple Lie group. Since $L\cong\mathbb{R}$, we may view $\Psi\colon L\to\operatorname{Aut}(N)$ as a one-parameter subgroup $s\mapsto\Psi(s)$. Note that $\operatorname{Aut}(N)=\operatorname{Aut}(\mathbb{T}^{k})\times\operatorname{Aut}(S)$, because every automorphism of $N$ must preserve the identity component of the center $Z^{0}(N)=\mathbb{T}^{k}$. Since $\operatorname{Aut}(\mathbb{T}^{k})=\operatorname{GL}(k,\mathbb{Z})$ is discrete, $\Psi(s)\in\operatorname{Aut}^{0}(S)$ for all $s\in\mathbb{R}$ (extended trivially to $N$). Note that because $S$ is compact and connected, by [23, Cor 4.48], the exponential map $\exp\colon\mathfrak{s}\to S$ is surjective. Therefore, the automorphism $\Psi(s)\in\operatorname{Aut}(S)$ is uniquely determined by $d\Psi(s)\in\operatorname{Aut}(\mathfrak{s})$. Then $s\mapsto d\Psi(s)$ is a one parameter subgroup of $\operatorname{Aut}(\mathfrak{s})$. By [23, Chap. I.14], the Lie algebra of $\operatorname{Aut}(\mathfrak{s})$ is $\operatorname{Der}(\mathfrak{s})=\operatorname{ad}(\mathfrak{s})$. Therefore, there exists a $V\in\mathfrak{s}$ such that
@@
It follows by the above and [23, Prop. 1.91] that $d\Psi(s)=e^{\operatorname{ad}(sV)}=\operatorname{Ad}(\exp(sV))$, where $e^{(\cdot)}\colon\operatorname{Der}(\mathfrak{s})\to\operatorname{Aut}(\mathfrak{s})$ is the exponential map, and hence
@@
It remains to deal with the case of general $N$. We do this by passing to a covering of the form dealt with in the previous case. By [23, Prop. 1.124] and [23, Cor. 4.25], the Lie algebra $\mathfrak{g}$ of $G$ is of the form
@@
where $\psi$ is the differential of $\Psi$, and $\mathfrak{s}$ is a compact semi-simple Lie algebra. Therefore, there is a compact, simply connected, semi-simple Lie group $S$ with Lie algebra $\mathfrak{s}$. Then $\mathbb{R}^{k}\times S$ has Lie algebra $\mathbb{R}^{k}\oplus\mathfrak{s}$. By [23, Thm. 1.125], there exists a unique map $\Phi\colon L\to\operatorname{Aut}(\mathbb{R}^{k}\times S)$ such that